OpenStudy (anonymous):

Find the Maclaurin series for f(x) = sin(x^4). Help...I'm completely stuck and confused :/

zepdrix (zepdrix):

Are we allowed to use the general form for the Maclaurin series and just plug stuff in?\[\Large\rm \sum_{0}^{\infty} \frac{f^{(n)}(0)}{n!}x^n\]

OpenStudy (anonymous):

Yes. :)

zepdrix (zepdrix):

So I guess calculating derivatives is going to be a little bit of a pain due to the chain rule :( Hmm.. I wonder if we can just take the sin(x) series and plug x^4 into it or something.... seems like it should work >.<

OpenStudy (anonymous):

Yeah..Well I have this so far... f(x) = sin(x^2); f'(x) = 2*cos(x^2)*x f''(x) = -4*sin(x^2)*x^2+2*cos(x^2) f'''(x) = -8*cos(x^2)*x^3-12*sin(x^2)*x f^(4)(x) = 16*sin(x^2)*x^4-48*cos(x^2)*x^2-12*sin(x^2)

OpenStudy (anonymous):

I think that's right if I did the derivatives correctly.

zepdrix (zepdrix):

Are you allowed to cheat and use the series for sin(x)? :O It works out REALLY nice and simple hehe.

zepdrix (zepdrix):

Wait why is it sin(x^2)? I thought it was sin(x^4), was that a typo?

OpenStudy (anonymous):

Ugh oops -_- Didn't think and put 2...Wow...xD

zepdrix (zepdrix):

heh

OpenStudy (anonymous):

Gotta start over.

OpenStudy (anonymous):

I don't know how I would set it up though...Once I get it

zepdrix (zepdrix):

If you insist on doing it the long way, I dunno lemme crunch out some derivatives also and see what we come up with.

zepdrix (zepdrix):

Little tip: if you're writing this all down on paper, use shorthand for sine and cosine, otherwise it gets really long and messy. Like I would write f(x)=Sx^4 or something

zepdrix (zepdrix):

Since it's a Maclaurin series, we'll need to evaluate each derivative at x=0. The problem you'll find is....... The x's don't start to "disappear" from in front of the trig functions until the 5th derivative or so. So you'll need to calculate to the 7th derivative or something like that... because the first 4 or so will all be zero :(

zepdrix (zepdrix):

Can we just do the shortcut way? >.<

OpenStudy (anonymous):

Whats the shortcut way? :)

zepdrix (zepdrix):

We can calculate the Maclaurin series for sin(x), (Taking a few derivatives and such if we need to) Eventually we'll end up with:\[\Large\rm \sin(x)=x+\frac{x^3}{3!}-\frac{x^5}{5!}+...\]And from there we just plug in x^4 for all the x's.

OpenStudy (anonymous):

Really? We can do that? :O

OpenStudy (anonymous):

Thats magical xD Hahah I never thought that was like "allowed" xP

zepdrix (zepdrix):

I wasn't sure either >.< I had to look it up. Seems to be legit though.

OpenStudy (anonymous):

Haha this will be fun. xD

OpenStudy (anonymous):

Well Im studying for a quiz and this was an example problem. It said to use the series I know.

zepdrix (zepdrix):

Ah very cool.

zepdrix (zepdrix):

What class?

OpenStudy (anonymous):

Physics class, doing some calculus problems to test us.

OpenStudy (anonymous):

Like a refresher

zepdrix (zepdrix):

Oh ok :O I dunno what field of study you're going into, I'm a math major myself. But if you're going to take any advanced math classes... make sure you have the power series for sin(x), cos(x) and even e^x TATTOOED into your brain. They show up everywhere.

OpenStudy (anonymous):

Thanks haha..Yeah I'm going into computer software and mechanical engineering. I will make sure to get them all down! So..The Maclaurin series expansion for sin(x) is \sin(x) =\sum_{k=0}^{\infty} \frac{(-1)^k}{(2k+1)!}x^{2k+1} = x -\frac{x^3}{3!}+\frac{x^5}{5!} - \ldots and the formula is valid for all real values of x.

OpenStudy (anonymous):

Ugh I cant at all use that dumb equation editor xD

zepdrix (zepdrix):

lol >.<

OpenStudy (anonymous):

So what happens if we replace all the x's with x^4?

OpenStudy (anonymous):

I mean...Sinx can be thought of as sinx^1...So I guess it makes sense to get rid of the trouble and think of it like that.

zepdrix (zepdrix):

Here is one of the series,\[\Large\rm \sum_{n=0}^{\infty}\frac{(-1)^n x^{2n}}{(2n)!}\] Here is the other trig function,\[\Large\rm \sum_{n=0}^{\infty}\frac{(-1)^n x^{2n+1}}{(2n+1)!}\] Here is how I'm able to remember which is which... Sin(x) is an `odd function`. Cos(x) is an `even function`. Can you tell which power series matches up with which using that info? ;)

OpenStudy (anonymous):

Sin(x) does :P

OpenStudy (anonymous):

I mean...Its not rocket science. sin x turns into cosx after taking the derivative.

zepdrix (zepdrix):

Maybe I made that confusing >.< I posted the series for sin(x) and cos(x) above. I was trying to see if you could tell which is which :P

OpenStudy (anonymous):

Ohhh xD Hahahah Yeah...I miss understood

OpenStudy (anonymous):

Sin(x) is odd...Cos(x) is even.

geerky42 (geerky42):

It is known that \(\sin(x) = x - \dfrac{x^3}{3!} + \dfrac{x^5}{5!} - \dfrac{x^7}{7!}+\ldots\) So you can easily replace \(x\) to \(x^4\) and you will get \(\sin(x^4) = (x^4) - \dfrac{(x^4)^3}{3!} + \dfrac{(x^4)^5}{5!} - \ldots = x^4 - \dfrac{x^{12}}{3!} + \dfrac{x^{20}}{5!}-\ldots\). Hence \[\sin(\color{red}{x}) = \sum_{k=0}^\infty \dfrac{\color{red}{x}^{2k+1}}{2k+1} \\ \sin(\color{blue}{x^4}) = \sum_{k=0}^\infty \dfrac{\color{blue}{x^4}^{2k+1}}{2k+1} = \boxed{\sum_{k=0}^\infty \dfrac{x^{8k+4}}{2k+1}}\]

geerky42 (geerky42):

Oops forgot factorials... haha

geerky42 (geerky42):

and (-1)^k... sorry for careless mistakes... but is it clear, right?

OpenStudy (anonymous):

Where does the (-1)^k come in? And that does make a lot of sense! THANK YOU!

geerky42 (geerky42):

Well, (-1)^k have to come in because sign changes from term to term

geerky42 (geerky42):

\(\displaystyle \sin x = \sum_{k=0}^\infty (-1)^k\dfrac{x^{2k+1}}{(2k+1)!}\)

OpenStudy (anonymous):

So its the same thing when we write sin(x^4)? It will have the answer you wrote but with (-1)^k

geerky42 (geerky42):

Of course. and (2k+1)! in denominator instead of just 2k+1 :)